Harlow Corp. uses a FIFO process costing system. Beginning inventory for January consisted of 1,000 units that were 60% completed. During January, 9,500 units were started and completed. On January 31, the inventory consisted of 500 units that were 70% completed. How many units were started during January?

a) 9,850

b) 10,000

c) 10,250

d) 11,000
Answer #1

b) 10,000

Units completed = 9,500 + 1,000 = 10,500

Units started = 10,500 + 500 - 1,000 = 10,000